Q1) Haroon’s budget line relating good X and good Y has intercept of 100 unit of good X and 40 units of good Y. if the price of good X is 24, what is Haroon’s income? What is the price of good Y? What is then slope of budget line?
1
Expert's answer
2020-06-15T11:12:18-0400

The budget line will take the form of:



"M = xP_x + yP_y"

Solving for "y" in the budget line:



"y = \dfrac{M}{P_y} - \dfrac{P_x}{P_y} x"

Where:


  • M is the income
  • "P_y" is the price of good Y and,
  • "P_x" is the price of good X.

Here, the slope of the budget line is "-\dfrac{P_x}{P_y}"

At the y-intercept, the value of good X consumed is zero. Therefore:



"M = yP_y"

Solving for ''y'', we get:



"y = \dfrac{M}{P_y}"

At the y-intercept, y = 40. Therefore:



"40 = \dfrac{M}{P_y}"

Solving for "P_y" :



"P_y = \dfrac{M}{40}"

At the x-intercept, the amount of Y consumed is zero. Therefore:



"M = xP_x"

Solving for ''x'':



"x = \dfrac{M}{P_x }"

The x-intercept is 100. Therefore:




"100 = \dfrac{M}{P_x }"

Solving for "P_x" :



"P_x = \dfrac{M}{100}"

Therefore, the slope is:



"-\dfrac{P_x}{P_y} = -\dfrac{\dfrac{M}{100}}{\dfrac{M}{40}}"

"-\dfrac{P_x}{P_y} = -\dfrac{M}{100}\times \dfrac{40}{M}"

"-\dfrac{P_x}{P_y} = \color{red}{-0.4}"